celticcat:

Statistics is not my strong point. How do you do this one:- The marks obtained by the class in a test are normally distributed with mean 74 and standard deviation 15. The top 10% get an A grade. Calculate the minimum mark required to get A grade.

celticcat:

I guess you need to use the standardised score equation z = (X - 74) / 15 but how do you get the z score?

Zowie:

z would be 0.1 because it's 10% but you have to change it to a decimal. When changing a percent to a decimal you move the . over 2 places to the right.

celticcat:

But that gives an answer of X = 75.5 but the correct answer is 94.

celticcat:

we have to use the Normal Distribution tables I guess...

celticcat:

Maybe I should be looking at the area under the Normal Cheeseulative curve of > 0.9???

celticcat:

What???? The computer isnt printing what Im typing!!!

Zowie:

It is. I just have no clue what you are talking about. Wouldn't want to give you a bad/wrong answer.

celticcat:

Yes thats fine. I meant to type C u m u l a t i ve.

Zarkon:

Solve \(\displaystyle\Phi^{-1}(.90)=\frac{x-74}{15}\) for \(x\) where \(\displaystyle\Phi^{-1}(\cdot)\) is the inverse standard normal CDF.

celticcat:

Ah right Than you.